OpenStudy (anonymous):

A 7-foot ladder is leaned against a building in such a way that the bottom of the ladder is 4 feet from the base of the wall. Find the angle formed between the ladder and the ground.

OpenStudy (anonymous):

Using the cos: cos (angle formed) = (Adjacent Side) / (Hypotenuse side) Adjacent Side = Distance from ladder to wall = 4 Hypotenuse side = Length of ladder The answer will be...

OpenStudy (anonymous):



OpenStudy (anonymous):

4/7 = What?

OpenStudy (anonymous):

@pcarr114 ?

OpenStudy (anonymous):

.5714285714

OpenStudy (anonymous):

OK now, @dauspex Shouldn't he find cos .571?

OpenStudy (anonymous):

I'm asking him because I might be wrong. :P

OpenStudy (anonymous):

but you are going the right way. :)

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

cos (angle) = 4 / 7 (angle) = cosh (4/7) * cosh = cos^-1